OpenStudy (abbycross167):

Can someone please help me with this math question? I will give a medal and will fan... Question - Divide (16x^6 – 12x^4+ 4x^2) by 4x^2

OpenStudy (zakariaars):

(4x^4)-(3x^2)+1 is what I got

OpenStudy (zakariaars):

First you divide each term by the 4x^2 16x^6/4x^2 would become 4(16/4)x^4(when dividing exponents, subtract them) Apply that to each term

OpenStudy (abbycross167):

16x^6/4x^2 = 4x^4 12x^4/4x^4 = 3x^2 4x^2/4x^4 = 1